INFRARED DEPLETION SPECTROSCOPY OF ANILINE-WATER_{n}(n=2-6) CATION CLUSTERS

\underline{T. NAKANAGA} and F. ITO, National Institute of Advanced Industrial Science and Technology, Higashi 1-1,Tsukuba, Ibaraki, 305-8565, Japan.

The vibrational spectra of aniline-water_{n} (n=2-6) cluster cations in the OH stretching vibration region have been measured by infrared depletion technique. The cluster cations were prepared by the near resonant two-photon ionization of neutral aniline-watern clusters. No structure was observed in the UV-excitation spectra, and the origin of the cluster cation seems to be the larger neutral aniline-water clusters. The observed mass spectrum showed magic numbers at n=4 and 5. The infrared spectrum of aniline-water2 shows strong absorption bands around 3200-3300cm^{−1}, and the absorption of the stretching band of free NH bond was observed at 3440cm^{−1}. This results suggests that one of the NH bonds of aniline interacts with the water dimer in the cluster cation. In the infrared spectra of larger aniline-water_{n}(n > 2) cations, strong absorption bands were observed at around 3400-3550cm^{−1}. The spectra resemble to those observed in the phenol-watern+ cation. This result suggests that the ring structure of water molecules interacting with the NH bonds of aniline cation.